OpenStudy (anonymous):

The x co-ordinate of a point on the curve 9y^2 = x^3 , the normal at which cuts off equal intercept on the axes is ?

OpenStudy (anonymous):

9y^2=x^3 d(9y^2)/dx =d(x^3)/dx 18y dy/dx =3x^2 dy/dx =x^2/(6y)

OpenStudy (anonymous):

The gradient of tangent is x^2/(6y) So, gradient of normal is -6y/(x^2)

OpenStudy (anonymous):

NOW, let x-intercept of the normal be a then y-intercept is also a.

OpenStudy (anonymous):

Gradient of normal = (a-0)/(0-a) -6y/(x^2) =-1

OpenStudy (anonymous):

Now, and solve for x

OpenStudy (anonymous):